A 65.0 mL sample of 1.0 M NaOH is mixed with 48.0 mL of 1.0 M H2SO4 in a large Styrofoam coffee cup; the cup is fitted with a lid through which passes a calibrated thermometer. The temperature of each solution before mixing is 24.4°C. After adding the NaOH solution to the coffee cup, the mixed solutions are stirred until reaction is complete. Assume that the density of the mixed solutions is 1.0 g/mL, that the specific heat of the mixed solutions is 4.18 J/(g •°C), and that no heat is lost to the surroundings. The ΔΗrxn for the neutralization of NaOH with H2SO4 is –114 kJ/mol H2SO4.

What is the maximum measured temperature in the Styrofoam cup?
Answer #1

65.0 mL sample of 1.0 M NaOH = 0.065 L * 1.0 mole / L = 0.065 mole

and

48.0 mL of 1.0 M H2SO4 = 0.048 L * 1.0 mole / L = 0.048 mole.

H2SO4 + 2 NaOH ...........> Na2SO4 + 2 H2O

one mole H2SO4 is neutralized by 2 mole NaOH.

NaOH is limiting reactant.

thus

mole of H2SO4 neutralized = 0.065 / 2 = 0.0325 mole.

thus

heat produced = 114 KJ / mole * 0.0325 mole = 3.705 KJ = 3705 J

total volume of solution = (48.0 + 65.0) = 113 ml

mass of solution mixture = 113 g

heat = m * s * dT

or

3705 = 113 * 4.18 * (T - 24.4)

or

T = 32.2 oC

thus

maximum measured temperature in the Styrofoam cup is 32.2 oC